OpenStudy (anonymous):

What is the ratio of a to b? 9 over 17= 3a over a +b????

OpenStudy (astrophysics):

\[\frac{ 9 }{ 17 } = \frac{ 3a }{ a+b }\]

OpenStudy (astrophysics):

You can cross multiply if you like, to get \[9(a+b)=(3)(17)a\]

OpenStudy (astrophysics):

You have to solve for a/b, so it will be fun distribute the 9 over (a+b) and what's 3*17